# BanglaTalk: Towards Real-Time Speech Assistance for Bengali Regional Dialects

### v1

@@ 1 Introduction
Conversational speech assistants (Dutsinma et al., 2022) have transformed human-computer interaction, making information more accessible. Widely adopted tools such as Alexa, Siri, and Cortana demonstrate the profound impact of real-time speech assistants on human lives (Hoy, 2018). However, while significant progress has been made for high-resource languages such as English, Mandarin, and French, such tools are still underdeveloped for the low-resource Bengali language. Bengali is a morphologically rich Indo-Aryan language (Islam et al., 2025), spoken by approximately 260 million people worldwide. It exhibits significant regional dialectal diversity, with variations in phonology, vocabulary, and syntax (Hasan et al., 2024b). This linguistic diversity poses a major challenge in building robust speech assistants.
@@
Automatic Speech Recognition (ASR) is a key component of speech assistant systems. Existing ASR systems are developed primarily for standard Bengali (Saha et al., 2021; Rakib et al., 2023b), and their performance is significantly degraded in regional dialects. As a result, existing speech assistant systems that integrate such ASR cannot support regional dialectal communication (Hasan et al., 2021; Arnab et al., 2023). Moreover, real-time deployment requires not only dialectal robustness, but also minimal end-to-end delay and efficient bandwidth usage. Previous works lack dialect-aware ASR, systematic analysis of delay minimization techniques, bandwidth efficiency, and real-time communication.
@@
In this work, we introduce BanglaTalk, the first real-time conversational speech assistant for Bengali regional dialects. BanglaTalk adopts a client-server architecture and incorporates the Real-time Transport Protocol (RTP) (Schulzrinne et al., 2003) to achieve low-latency communication. Robust audio encoding enables operation at 24 kbps (kilobits per second). As illustrated in Fig.˜1, while existing speech assistants fail in interpreting regional dialectal queries, BanglaTalk transcribes them accurately through the dialect-aware ASR system. It responds to queries effectively and interactively in real-time using low bandwidth.
@@
The BanglaTalk client integrates lightweight audio processing modules, including noise cancellation, dynamic range compression, and audio encoding. On the server side, a dialect-aware ASR system, a voice activity detector (VAD), a natural-sounding Text-to-Speech (TTS) system, and audio encoding modules form a complete pipeline for real-time speech assistance. Central to this system is BRDialect, a dialect-aware ASR model fine-tuned on ten Bengali regional dialects. BRDialect outperforms the baseline Whisper (Tugstugi, 2023) and IndicWav2Vec (Javed et al., 2022) models, achieving a word error rate of 74.1% and character error rate of 40.6% on the RegSpeech12 (Hassan et al., 2025) dataset.
@@
Additionally, the integrated VITS (Kim et al., 2021) TTS model produces natural-sounding speech with a high mean opinion score (MOS) of 4.49, enhancing the user experience. With an average end-to-end delay of 4.9 seconds, BanglaTalk enables interactive real-time communication between the user and the speech assistant. This system will significantly impact the lives of Bengali speakers due to its dialect-aware ASR, low bandwidth usage, and real-time performance.
@@ 2.1 BanglaTalk Client
The client captures audio in 20-ms (milliseconds) frames at a sample rate of 16 kHz (kilohertz). Each frame contains 320 samples in 16-bit PCM little-endian format (Dobson, 2000). Although the Opus codec (Valin et al., 2016) supports multiple sample rates (e.g., 8-48 kHz), we fix the sample rate at 16 kHz to align with the ASR and VAD modules. Opus allows frame durations of 2.5-100 ms. A frame duration of 20 ms (50 RTP packets per second) offers a balance between packet size and loss rate in real-time communication.
@@
Speech captured from the microphone often includes soft and excessively loud signals. Compressing the dynamic range helps maintain a consistent audio level, enhancing performance (Giannoulis et al., 2012). We develop the dynamic range compression algorithm described in Alg.˜1 and apply compression only to the loud audio segments. Whenever the decibel level of a normalized audio sample exceeds -10 dBFS (decibels relative to full scale), we apply a compression ratio of 2:1. Samples outside this threshold remain unchanged.
@@
Noise poses a major issue in audio communication. Background and foreground noise can be picked up by the microphone and transmitted to the network, degrading the overall performance of the system. To mitigate this, we perform noise suppression with RNNoise (Valin, 2018). It is a lightweight neural network-based denoiser capable of real-time operation.
RNNoise is trained to remove noise from an audio frame of duration 10 ms at a 48 kHz sample rate. Since every captured audio frame duration is 20 ms at a 16 kHz sample rate, we apply audio segmentation, upsampling, and downsampling during noise suppression. Alg.˜2 shows the pseudocode for upsampling. Linear interpolation (Xu and Xu, 2022) is used in upsampling from 16 to 48 kHz due to minimal computational overhead. After denoising, the audio is downsampled to 16 kHz using Alg.˜3. Specifically, we skip intermediate sample values – from every three consecutive audio samples, the first one is retained and the remaining two are discarded. This simple technique is computationally efficient for downsampling.
@@
Opus is a high-quality audio codec for interactive speech and music transmission over the Internet (Valin et al., 2016). It is widely used in VoIP (Voice over IP) applications (Sundvall, 2014) due to the low latency processing and error concealment. Each audio frame contains a total of 320 samples (640 bytes). Without compression, audio frames are expensive to transmit over the Internet due to high bandwidth consumption. To mitigate this, each audio frame is encoded using the Opus codec with a low bitrate of 24 kbps. This ensures low bandwidth usage, which is a critical factor for greater accessibility. The incoming audio frames from the server are decoded with the same codec.
@@
Each encoded audio frame is encapsulated in an RTP packet following the Real-time Transport Protocol (Group et al., 1996). The first 12 bytes of each packet contain the header, and the remaining bytes contain the Opus-encoded payload. App.˜B describes the structure of the RTP packet. These packets are sent at 20 ms intervals to the server’s public IP and the RTP packet receiver port (Postel, 1980).
@@ 2.2 BanglaTalk Server
Voice activity detection is a crucial component of real-time communication, as it identifies speech segments while discarding non-speech portions. This prevents unnecessary downstream processing, thus reducing latency and improving overall efficiency. We use the Silero VAD (Team, 2024) in the streaming mode, which is specifically designed for real-time applications. It works with a frame duration of 32 ms (512 samples) at a sample rate of 16 kHz and determines whether each audio frame marks the beginning, end of a speech segment, or none. Only the detected speech segment corresponding to the user query is forwarded to the ASR system.
@@
We train a Wav2Vec2-based model (Baevski et al., 2020) using speech data from ten regional dialects from the Ben10 dataset (Humayun et al., 2024). For data processing, we follow Hasan et al. (2024a) and fine-tune the pre-trained IndicWav2Vec (Javed et al., 2022) for Bengali on that processed dataset. To evaluate the performance of our trained ASR model, we use the RegSpeech12 (Hassan et al., 2025) test set (Ben10 test set is not publicly available). A detailed description and analysis of these datasets are provided in §˜C.1 and §˜C.2.
@@
Accurate detection of the end of a user query is essential for real-time speech assistance systems (Liang et al., 2023). We have defined the end of query as a silence segment lasting at least 1.2 seconds. For silence detection, Silero VAD is utilized. Once an end-of-query is detected, the speech segment is passed to the ASR system to generate the transcription. The resulting query is then forwarded to the LLM, which produces the system’s response.
@@
Large Language Models (LLMs) are crucial for generating responses to user queries (Dam et al., 2024). To maintain coherent communication, responses are generated for valid queries, while invalid queries are discarded. We employ GPT-4.1-nano as the chat model, with the prompt template presented in App.˜D. To minimize latency, we use streaming mode, which delivers responses incrementally rather than waiting for a full response. The streamed text is segmented based on Bengali punctuation and forwarded to the TTS system.
@@
Several TTS models are available for Bengali (Raju et al., 2019). We experiment with MMS-TTS-Ben (Pratap et al., 2024) and two variants of VITS-Bengali (male and female voices) (Hossen, 2023). These models are selected because of their minimal processing delay. MMS-TTS-Ben produces speech at a 16 kHz sample rate, while VITS-Bengali outputs at 22.05 kHz. For system compatibility, the VITS-generated speech is resampled to 16 kHz.
@@ 3.2 Evaluating BRDialect
To evaluate the performance of our dialect-aware ASR system, BRDialect, we use the test set from the RegSpeech12 (Hassan et al., 2025) dataset. It includes dialects from twelve regions of Bangladesh – Rangpur, Sylhet, Chittagong, Noakhali, Narail, Kishoreganj, Barishal, Habiganj, Comilla, Tangail, Sandwip, and Narsingdi, totaling 2132 audio files.
For evaluation, we report the Word Error Rate (WER) and Character Error Rate (CER), following the formulas described in §˜E.1. To refine ASR predictions, we apply beam search decoding with a 5-gram KenLM language model (Heafield, 2011). We also investigated the impact of the preprocessing and postprocessing steps, including noise cancellation, normalization, and punctuation removal, as these factors significantly affect the overall performance of the ASR.
@@
BRDialect consistently outperforms Whisper-medium-Bengali (Tugstugi, 2023) and IndicWav2Vec-Bengali (Javed et al., 2022), achieving a 12.41–17.39% relative improvement in WER and a 27.77–33.98% improvement in CER. The comparatively poor performance of the baseline models suggests that dialectal variation is not adequately addressed during their training, limiting their suitability for regional speech-to-text tasks. BRDialect highlights the importance of dialect-aware fine-tuning for building a robust ASR system.
@@
We further evaluate BRDialect across individual regions. As shown in Fig.˜5, the ASR system performs well across most regions, with WER below 70% in seven out of the twelve regions. The lowest WER, 0.438, is achieved for the Comilla region.
@@
5-gram KenLM Decoding
Integrating a 5-gram KenLM model during decoding improves WER from 0.865 to 0.827, a 4.39% reduction. This confirms the value of training a KenLM language model with a Bengali regional text corpus for robust ASR performance (Rakib et al., 2023b).
Unicode Normalization
Normalizing text further reduces WER to 0.796. This step is crucial since many Bengali characters can be represented in multiple ways, causing transcription inconsistencies. Normalizing text with the BnUnicodeNormalizer (Ansary et al., 2023) ensures uniform representation, improving ASR performance.
Punctuation Removal
@@
Levenshtein Distance Analysis
Fig.˜3 illustrates the distribution of normalized Levenshtein distance between the ground-truth transcriptions and the outputs of the evaluated ASR systems under the best processing configuration – No noise cancellation, Unicode normalization, and punctuation removal. For BRDialect, transcription quality is further refined during decoding using our trained 5-gram KenLM language model. A lower Levenshtein distance indicates higher transcription accuracy.
@@ 3.3 Evaluating BanglaTalk
Among these, the encoding with the Opus codec is the most efficient, requiring only 0.56 ms per frame. This is consistent with the codec’s design, which targets real-time, low-latency audio applications (Valin et al., 2016). Similarly, the DRC module introduces minimal computational overhead, averaging 1.31 ms per frame. In contrast, the noise cancellation with RNNoise introduces the highest processing time of 6.51 ms.
RNNoise applies a lightweight neural network for speech enhancement, making it more resource-intensive than traditional signal processing methods. Additionally, RNNoise operates exclusively at a 48 kHz sample rate, while the BanglaTalk system relies on 16 kHz to ensure compatibility with ASR and VAD modules (see §˜2.1). As a result, each frame must first be upsampled from 16 kHz to 48 kHz prior to noise cancellation and subsequently downsampled back to 16 kHz. These resampling operations introduce an additional delay of 1.40 ms for upsampling and 0.34 ms for downsampling. Compared to other ML-based noise cancellation systems (Cha et al., 2023), RNNoise provides an excellent balance of speech quality and efficiency, making it highly suitable for real-time speech applications (Valin, 2018).
@@
In real-time assistive speech technologies, both end-to-end latency and bandwidth efficiency are crucial factors (Kaur et al., 2019). Bandwidth efficiency is particularly important for underserved populations, such as those in rural areas of Bangladesh, where high-speed or expensive Internet connections are often unavailable.
To address this challenge, our system prioritizes bandwidth efficiency without compromising performance. Audio streams are encoded at a target bitrate of 24 kbps, significantly reducing bandwidth usage in the upload and download streams during client-server communication. Fig.˜4 shows the upload bitrate of client-side for a one-minute audio stream.
We employ the variable bitrate (VBR) mode of the Opus codec, which dynamically adjusts the bitrate based on the characteristics of the audio signal – allocating more bitrate to speech segments and conserving bitrate during silence. In the example shown in Fig.˜4, the average bitrate over one minute is only 19.29 kbps. This adaptive bitrate usage makes the system more accessible to users in economically disadvantaged regions (Osuagwu et al., 2013), who might otherwise be unable to benefit from assistive speech technologies.
@@
We conduct rigorous testing and time analysis to measure the BanglaTalk system’s end-to-end delay. As detailed in §˜E.3, the BRDialect ASR module of the BanglaTalk system introduces only a small delay, making it well-suited for real-time applications. The TTS systems evaluated in our study also exhibit low processing delay. App.˜F presents a detailed evaluation of TTS systems. Among them, the VITS-Bengali (male variant) is the best-performing model, achieving a high Mean Opinion Score (MOS) of 4.49 on the subset of the BanSpeech (Samin et al., 2024) dataset.

### v2

@@ 1 Introduction
Conversational speech assistants (dutsinma2022systematic) have transformed human-computer interaction, making information more accessible. Widely adopted tools such as Alexa, Siri, and Cortana demonstrate the profound impact of real-time speech assistants on human lives (hoy2018alexa). However, while significant progress has been made for high-resource languages such as English, Mandarin, and French, such tools are still underdeveloped for the low-resource Bengali language. Bengali is a morphologically rich Indo-Aryan language (islam2025banglalem), spoken by approximately 260 million people worldwide. It exhibits significant regional dialectal diversity, with variations in phonology, vocabulary, and syntax (hasan2024credibility). This linguistic diversity poses a major challenge in building robust speech assistants.
@@
Automatic Speech Recognition (ASR) is a key component of speech assistant systems. Existing ASR systems are developed primarily for standard Bengali (saha2021development; rakib2023bangla), and their performance is significantly degraded in regional dialects. As a result, existing speech assistant systems that integrate such ASR cannot support regional dialectal communication (hasan2021alapi; arnab2023shohojogi). Moreover, real-time deployment requires not only dialectal robustness, but also minimal end-to-end delay and efficient bandwidth usage. Previous works lack dialect-aware ASR, systematic analysis of delay minimization techniques, bandwidth efficiency, and real-time communication.
@@
In this work, we introduce BanglaTalk, the first real-time conversational speech assistant for Bengali regional dialects. BanglaTalk adopts a client-server architecture and incorporates the Real-time Transport Protocol (RTP) (schulzrinne2003rtp) to achieve low-latency communication. Robust audio encoding enables operation at 24 kbps (kilobits per second). As illustrated in Fig.˜1, while existing speech assistants fail in interpreting regional dialectal queries, BanglaTalk transcribes them accurately through the dialect-aware ASR system. It responds to queries effectively and interactively in real-time using low bandwidth.
@@
The BanglaTalk client integrates lightweight audio processing modules, including noise cancellation, dynamic range compression, and audio encoding. On the server side, a dialect-aware ASR system, a voice activity detector (VAD), a natural-sounding Text-to-Speech (TTS) system, and audio encoding modules form a complete pipeline for real-time speech assistance. Central to this system is BRDialect, a dialect-aware ASR model fine-tuned on ten Bengali regional dialects. BRDialect outperforms the baseline Whisper (BengaliAIASRSubmission2023) and IndicWav2Vec (javed2022towards) models, achieving a word error rate of 74.1% and character error rate of 40.6% on the RegSpeech12 (regspeech12) dataset.
@@
Additionally, the integrated VITS (kim2021conditional) TTS model produces natural-sounding speech with a high mean opinion score (MOS) of 4.49, enhancing the user experience. With an average end-to-end delay of 4.9 seconds, BanglaTalk enables interactive real-time communication between the user and the speech assistant. This system will significantly impact the lives of Bengali speakers due to its dialect-aware ASR, low bandwidth usage, and real-time performance.
@@ 2.1 BanglaTalk Client
The client captures audio in 20-ms (milliseconds) frames at a sample rate of 16 kHz (kilohertz). Each frame contains 320 samples in 16-bit PCM little-endian format (dobson2000developments). Although the Opus codec (valin2016high) supports multiple sample rates (e.g., 8-48 kHz), we fix the sample rate at 16 kHz to align with the ASR and VAD modules. Opus allows frame durations of 2.5-100 ms. A frame duration of 20 ms (50 RTP packets per second) offers a balance between packet size and loss rate in real-time communication.
@@
Speech captured from the microphone often includes soft and excessively loud signals. Compressing the dynamic range helps maintain a consistent audio level, enhancing performance (giannoulis2012digital). We develop the dynamic range compression algorithm described in Alg.˜1 and apply compression only to the loud audio segments. Whenever the decibel level of a normalized audio sample exceeds -10 dBFS (decibels relative to full scale), we apply a compression ratio of 2:1. Samples outside this threshold remain unchanged.
@@
Noise poses a major issue in audio communication. Background and foreground noise can be picked up by the microphone and transmitted to the network, degrading the overall performance of the system. To mitigate this, we perform noise suppression with RNNoise (valin2018hybrid). It is a lightweight neural network-based denoiser capable of real-time operation.
RNNoise is trained to remove noise from an audio frame of duration 10 ms at a 48 kHz sample rate. Since every captured audio frame duration is 20 ms at a 16 kHz sample rate, we apply audio segmentation, upsampling, and downsampling during noise suppression. Alg.˜2 shows the pseudocode for upsampling. Linear interpolation (xu2022research) is used in upsampling from 16 to 48 kHz due to minimal computational overhead. After denoising, the audio is downsampled to 16 kHz using Alg.˜3. Specifically, we skip intermediate sample values – from every three consecutive audio samples, the first one is retained and the remaining two are discarded. This simple technique is computationally efficient for downsampling.
@@
Opus is a high-quality audio codec for interactive speech and music transmission over the Internet (valin2016high). It is widely used in VoIP (Voice over IP) applications (sundvall2014opus) due to the low latency processing and error concealment. Each audio frame contains a total of 320 samples (640 bytes). Without compression, audio frames are expensive to transmit over the Internet due to high bandwidth consumption. To mitigate this, each audio frame is encoded using the Opus codec with a low bitrate of 24 kbps. This ensures low bandwidth usage, which is a critical factor for greater accessibility. The incoming audio frames from the server are decoded with the same codec.
@@
Each encoded audio frame is encapsulated in an RTP packet following the Real-time Transport Protocol (audio1996rfc1889). The first 12 bytes of each packet contain the header, and the remaining bytes contain the Opus-encoded payload. App.˜B describes the structure of the RTP packet. These packets are sent at 20 ms intervals to the server’s public IP and the RTP packet receiver port (postel1980user).
@@ 2.2 BanglaTalk Server
Voice activity detection is a crucial component of real-time communication, as it identifies speech segments while discarding non-speech portions. This prevents unnecessary downstream processing, thus reducing latency and improving overall efficiency. We use the Silero VAD (SileroVAD) in the streaming mode, which is specifically designed for real-time applications. It works with a frame duration of 32 ms (512 samples) at a sample rate of 16 kHz and determines whether each audio frame marks the beginning, end of a speech segment, or none. Only the detected speech segment corresponding to the user query is forwarded to the ASR system.
@@
We train a Wav2Vec2-based model (baevski2020wav2vec) using speech data from ten regional dialects from the Ben10 dataset (ben10). For data processing, we follow hasan_saim_mostafa_2024_bengaliASR_indicwav2vec and fine-tune the pre-trained IndicWav2Vec (javed2022towards) for Bengali on that processed dataset. To evaluate the performance of our trained ASR model, we use the RegSpeech12 (regspeech12) test set (Ben10 test set is not publicly available). A detailed description and analysis of these datasets are provided in §˜C.1 and §˜C.2.
@@
Accurate detection of the end of a user query is essential for real-time speech assistance systems (liang2023dynamic). We have defined the end of query as a silence segment lasting at least 1.2 seconds. For silence detection, Silero VAD is utilized. Once an end-of-query is detected, the speech segment is passed to the ASR system to generate the transcription. The resulting query is then forwarded to the LLM, which produces the system’s response.
@@
Large Language Models (LLMs) are crucial for generating responses to user queries (dam2024complete). To maintain coherent communication, responses are generated for valid queries, while invalid queries are discarded. We employ GPT-4.1-nano as the chat model, with the prompt template presented in App.˜D. To minimize latency, we use streaming mode, which delivers responses incrementally rather than waiting for a full response. The streamed text is segmented based on Bengali punctuation and forwarded to the TTS system.
@@
Several TTS models are available for Bengali (raju2019bangla). We experiment with MMS-TTS-Ben (pratap2024scaling) and two variants of VITS-Bengali (male and female voices) (ComprehensiveBanglaTTS2023). These models are selected because of their minimal processing delay. MMS-TTS-Ben produces speech at a 16 kHz sample rate, while VITS-Bengali outputs at 22.05 kHz. For system compatibility, the VITS-generated speech is resampled to 16 kHz.
@@ 3.2 Evaluating BRDialect
To evaluate the performance of our dialect-aware ASR system, BRDialect, we use the test set from the RegSpeech12 (regspeech12) dataset. It includes dialects from twelve regions of Bangladesh – Rangpur, Sylhet, Chittagong, Noakhali, Narail, Kishoreganj, Barishal, Habiganj, Comilla, Tangail, Sandwip, and Narsingdi, totaling 2132 audio files.
For evaluation, we report the Word Error Rate (WER) and Character Error Rate (CER), following the formulas described in §˜E.1. To refine ASR predictions, we apply beam search decoding with a 5-gram KenLM language model (heafield2011kenlm). We also investigated the impact of the preprocessing and postprocessing steps, including noise cancellation, normalization, and punctuation removal, as these factors significantly affect the overall performance of the ASR.
@@
BRDialect consistently outperforms Whisper-medium-Bengali (BengaliAIASRSubmission2023) and IndicWav2Vec-Bengali (javed2022towards), achieving a 12.41–17.39% relative improvement in WER and a 27.77–33.98% improvement in CER. The comparatively poor performance of the baseline models suggests that dialectal variation is not adequately addressed during their training, limiting their suitability for regional speech-to-text tasks. BRDialect highlights the importance of dialect-aware fine-tuning for building a robust ASR system.
@@
We further evaluate BRDialect across individual regions. As shown in Fig.˜3, the ASR system performs well across most regions, with WER below 70% in seven out of the twelve regions. The lowest WER, 0.438, is achieved for the Comilla region. Although the Comilla dialect is not included in the training data of BRDialect, its low WER demonstrates the model’s strong generalization capability to unseen dialects.
@@ 3.2.1 Ablation Study
Impact of 5-gram KenLM Decoding
Integrating a 5-gram KenLM model during decoding improves WER from 0.865 to 0.827, a 4.39% reduction. This confirms the value of training a KenLM language model with a Bengali regional text corpus for robust ASR performance (rakib2023bangla).
Impact of Unicode Normalization
Normalizing text further reduces WER to 0.796. This step is crucial since many Bengali characters can be represented in multiple ways, causing transcription inconsistencies. Normalizing text with the BnUnicodeNormalizer (ansary2023unicode) ensures uniform representation, improving ASR performance.
Impact of Punctuation Removal
@@
Impact of High WER
A Word Error Rate (WER) of 74.1% is relatively high for a general-purpose ASR system. However, dialectal speech recognition is inherently more challenging than the standard ASR task. The poor performance of the baseline models (Whisper-medium-Bengali and IndicWav2Vec-Bengali) validates this difficulty. In this context, BRDialect achieves a relative improvement of 12.41-33.98% over the baseline models, representing a substantial achievement.
Within the BanglaTalk system, the integrated large language model (GPT-4.1-nano) effectively compensates for minor transcription errors. Leveraging its robust contextual understanding, the LLM can infer user intent even from noisy or imperfect ASR outputs, as illustrated in Fig.˜10. Most observed errors involve minor substitutions that do not significantly affect the intended meaning.
Furthermore, explicitly prompting the LLM to interpret dialectal queries (Fig.˜7) enhances the system’s ability to generate appropriate responses. The combination of a dialect-aware ASR model (BRDialect) and the powerful LLM (GPT-4.1-nano) ensures that the BanglaTalk system remains usable and effective despite relatively imperfect transcriptions.
3.2.2 Levenshtein Distance Analysis
Fig.˜4 illustrates the distribution of normalized Levenshtein distance between the ground-truth transcriptions and the outputs of the evaluated ASR systems under the best processing configuration – No noise cancellation, Unicode normalization, and punctuation removal. For BRDialect, transcription quality is further refined during decoding using our trained 5-gram KenLM language model. A lower Levenshtein distance indicates higher transcription accuracy.
@@ 3.3 Evaluating BanglaTalk
Among these, the encoding with the Opus codec is the most efficient, requiring only 0.56 ms per frame. This is consistent with the codec’s design, which targets real-time, low-latency audio applications (valin2016high). Similarly, the DRC module introduces minimal computational overhead, averaging 1.31 ms per frame. In contrast, the noise cancellation with RNNoise introduces the highest processing time of 6.51 ms.
RNNoise applies a lightweight neural network for speech enhancement, making it more resource-intensive than traditional signal processing methods. Additionally, RNNoise operates exclusively at a 48 kHz sample rate, while the BanglaTalk system relies on 16 kHz to ensure compatibility with ASR and VAD modules (see §˜2.1). As a result, each frame must first be upsampled from 16 kHz to 48 kHz prior to noise cancellation and subsequently downsampled back to 16 kHz. These resampling operations introduce an additional delay of 1.40 ms for upsampling and 0.34 ms for downsampling. Compared to other ML-based noise cancellation systems (cha2023dnoisenet), RNNoise provides an excellent balance of speech quality and efficiency, making it highly suitable for real-time speech applications (valin2018hybrid).
@@
In real-time assistive speech technologies, both end-to-end latency and bandwidth efficiency are crucial factors (kaur2019intelligent). Bandwidth efficiency is particularly important for underserved populations, such as those in rural areas of Bangladesh, where high-speed or expensive Internet connections are often unavailable.
To address this challenge, our system prioritizes bandwidth efficiency without compromising performance. Audio streams are encoded at a target bitrate of 24 kbps, significantly reducing bandwidth usage in the upload and download streams during client-server communication. Fig.˜5 shows the upload bitrate of client-side for a one-minute audio stream.
We employ the variable bitrate (VBR) mode of the Opus codec, which dynamically adjusts the bitrate based on the characteristics of the audio signal – allocating more bitrate to speech segments and conserving bitrate during silence. In the example shown in Fig.˜5, the average bitrate over one minute is only 19.29 kbps. This adaptive bitrate usage makes the system more accessible to users in economically disadvantaged regions (osuagwu2013low), who might otherwise be unable to benefit from assistive speech technologies.
Although applying RNNoise noise cancellation slightly increases the ASR system’s Word Error Rate (WER) by 1.27% (Table˜2), it is retained in the client application due to its substantial benefit in reducing bandwidth usage. By removing background and foreground noise, the audio signal becomes more compressible. Without noise cancellation, the average upload bitrate of the audio (as shown in Fig.˜5) rises from 19.29 kbps to 23.6 kbps – an 18.26% increase that is inefficient for client-side transmission. Given that the large language model (GPT-4.1-nano) demonstrates strong robustness in understanding minor transcription errors and generating accurate responses (Fig.˜10), the advantages of incorporating RNNoise into the client application outweigh its modest negative impact on ASR performance.
@@
We conduct rigorous testing and time analysis to measure the BanglaTalk system’s end-to-end delay. As detailed in §˜E.3, the BRDialect ASR module of the BanglaTalk system introduces only a small delay, making it well-suited for real-time applications. The TTS systems evaluated in our study also exhibit low processing delay. App.˜F presents a detailed evaluation of TTS systems. Among them, the VITS-Bengali (male variant) is the best-performing model, achieving a high Mean Opinion Score (MOS) of 4.49 on the subset of the BanSpeech (samin2024banspeech) dataset.
@@
Preliminary User Study
To evaluate the real-world usability of the BanglaTalk system, we conduct a limited user study involving four native speakers of the Sylhet and Mymensingh dialects. The Sylhet dialect is selected due to its relative high WER (0.831), while the Mymensingh dialect is included to assess the generalizability of BRDialect to unseen dialects, as it is not part of the training set. Two native speakers from each region interact with BanglaTalk on general information and everyday task queries. Participants rate their interaction experience on a 1-5 scale, where 1 indicates a poor experience and 5 indicates an excellent experience. Table˜9 summarizes the results of five queries per user. Overall, BanglaTalk achieves a mean rating of 3.62, indicating a generally positive user experience.
As illustrated in Fig.˜10, users from both regions interact successfully with the system despite minor transcription errors. Notably, users from the Mymensigh region, whose dialect is not included in BRDialect’s training data, report a mean satisfaction rating of 3.73, further demonstrating the model’s strong generalization capability to unseen dialects. Participants also note that, despite their strong accent in regional dialects, BanglaTalk responds accurately and naturally, providing an interactive and effective speech-based experience.